# Electron polarization induced by electroweak interaction in high-energy scattering off light nuclei

### v2

@@ Abstract
Overall, all the above ratios show a little difference among selected nuclei in the scattering processes at energies below 10 GeV. At higher energies, their magnitudes increase in the order of 7Li, 6Li, and 7Be nuclei. Particularly, there are remarkable discrepancies in the contribution of electron polarization to the scattering cross section when comparing stable and unstable nuclei. The 7Li -7Be pair proves this remark, although they have the same nucleon numbers. In contrast, for the 6Li -7Li stable pair, these discrepancies are much smaller, even though their nucleon numbers are slightly different. It can be explained that the weak interaction between electrons and nuclei strongly depends on the nuclear stability. This property affects the correlation between the weak interaction and electron polarization, along with their contribution in each scattering process.
Furthermore, although the $\sigma_{\text{total}}^{10}$ term can take negative values, its magnitude is significantly smaller than the $\sigma_{\text{total}}^{0}$ and $\sigma_{\text{total}}^{11}$ ones. Therefore, the total cross section also depends on the correlation between the electron spin orientations before and after scattering and remains positive in all cases. In particular, when the longitudinally polarized electrons move straight forward after scattering, the electron polarization-weak interaction correlation is not exhibited. This is evidenced by the ratios at $\theta=0^{\circ}$ in Tab. I. Their values remain nearly unchanged at all considered energies, except for $\sigma_{\text{total}}^{10}$, whose contribution can be omitted. So, besides nuclear spin orientation, it is evident that electron polarization and its transformation manner under the weak interaction play a crucial role in generating various types of asymmetry in electron-nucleus scattering at high energies.
@@
In conclusion, we have performed a systematic study on the impact of electron polarization on scattering off unoriented light nuclei using the unified electroweak theory. Numerical calculations of the scattering cross sections in different electron polarization cases for three 6,7Li and 7Be light nuclei are obtained by directly calculating their multipole form factors, using the formulas provided in Ref. [14]. The results obtained are then compared and/or correlated with the corresponding cross sections for the unpolarized electron scattering case. It has been found that at the scattering angle of $\theta=0^{\circ}$, the longitudinal electron polarization does not explicitly correlate with the weak interaction for all energy scales. Such a correlation is strongly exhibited at nonzero scattering angles ($\theta=5^{\circ}-8^{\circ}$) and high energies above 10 GeV. By comparing the results obtained for stable and unstable nuclei, we found that the nuclear stability significantly influences the contribution of electron polarization to the scattering cross section. This research direction can be extended in many ways, specifically considering various state transitions within a single nucleus or different nuclei. Then, the asymmetry resulting solely from electron polarization and the weak interaction can be figured out. In addition, other interesting results are expected once the nuclear spin orientation and charge-changing weak current are included. Moreover, the multipole expansion technique for the scattering cross section within the unified electroweak theory also suggests a new approach to directly study the EMC effect in a target nucleus using both unpolarized and polarized electrons. More specifically, we calculate the accumulative cross section contributing by separate nucleons inside a certain nucleus and compare it with that made of a similar number of nucleons bound inside the same one. It means that we can investigate the above effect upon only a single nucleus instead of indirectly comparing its scattering cross section with that of a lighter nucleus. These extensive studies will provide further vital information, contributing to a thorough understanding of the nature of the EMC effect in deep inelastic electron-nucleus scattering.
@@
Table 1: Calculated cross section ratios for 6,7Li and 7Be obtained within the unified electroweak theory at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and small scattering angles.
@@
10 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 9.8237 ×$\times 10^{-14}$ 4.6965 ×$\times 10^{-12}$ 2.1329 ×$\times 10^{-12}$ 9.8237 ×$\times 10^{-14}$ 4.6965 ×$\times 10^{-12}$ 2.1329 ×$\times 10^{-12}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 4.2043 ×$\times 10^{-5}$ 3.2235 ×$\times 10^{-5}$ 8.4485 ×$\times 10^{-6}$ 4.2047 ×$\times 10^{-5}$ 3.2237 ×$\times 10^{-5}$ 8.4486 ×$\times 10^{-6}$ 1.0001 1.0001 1.0000
$5^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0001 1.$0$2.4903 ×$\times 10^{-5}$ 8.9254 ×$\times 10^{-5}$1.4668 ×$\times 10^{-5}$2.4902 ×$\times 10^{-5}$ 8.9270 ×$\times 10^{-5}$1.4667 ×$\times 10^{-5}$ 1.0000 1.0002 1.0000
$8^{\circ}$ 0.4999 0.5001 0.4998 0.9998 1.0002 0.9997 -2.0072 ×$\times 10^{-4}$ 2.2865 ×$\times 10^{-4}$3.4970 ×$\times 10^{-4}$2.0063 ×$\times 10^{-4}$ 2.2876 ×$\times 10^{-4}$3.4946 ×$\times 10^{-4}$ 0.9996 1.0005 0.9993
100 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 9.8237 ×$\times 10^{-12}$ 4.6965 ×$\times 10^{-10}$ 2.1329 ×$\times 10^{-10}$ 9.8237 ×$\times 10^{-12}$ 4.6965 ×$\times 10^{-10}$ 2.1329 ×$\times 10^{-10}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5008 0.5016 0.5028 1.0016 1.0032 1.0055 0.0016 0.0032 0.0055 0.0016 0.0032 0.0055 1.0032 1.0064 1.0111
$5^{\circ}$ 0.5023 0.5044 0.5082 1.0045 1.0088 1.0165 0.0045 0.0087 0.0159 0.0045 0.0088 0.0165 1.0090 1.0177 1.0329
$8^{\circ}$ 0.5054 0.5112 0.5219 1.0109 1.0223 1.0438 0.0106 0.0214 0.0403 0.0109 0.0223 0.0438 1.0218 1.0447 1.0877
500 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 2.4559 ×$\times 10^{-10}$ 1.1741 ×$\times 10^{-8}$ 5.3315 ×$\times 10^{-9}$ 2.4559 ×$\times 10^{-10}$ 1.1741 ×$\times 10^{-8}$ 5.3315 ×$\times 10^{-9}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5173 0.5371 0.5753 1.0346 1.0741 1.1506 0.0324 0.0645 0.1157 0.0346 0.0741 0.1506 1.0692 1.1482 1.3011
$5^{\circ}$ 0.5414 0.5894 0.6793 1.0829 1.1788 1.3587 0.0711 0.1317 0.2088 0.0829 0.1788 0.3587 1.1657 1.3577 1.7173
$8^{\circ}$ 0.5811 0.6722 0.8217 1.1622 1.3444 1.6435 0.1225 0.2039 0.2814 0.1622 0.3444 0.6435 1.3244 1.6889 2.2870
1000 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000 9.8236 ×$\times 10^{-10}$ 4.6965 ×$\times 10^{-8}$ 2.1317 ×$\times 10^{-8}$ 9.8236 ×$\times 10^{-10}$ 4.6965 ×$\times 10^{-8}$ 2.1317 ×$\times 10^{-8}$ 1.0000 1.0000 1.0000
$3^{\circ}$ 0.5550 0.6183 0.7326 1.1100 1.2365 1.4652 0.0901 0.1606 0.2410 0.1100 0.2365 0.4652 1.2199 1.4730 1.9305
$5^{\circ}$ 0.6041 0.7173 0.8837 1.2081 1.4346 1.7674 0.1470 0.2325 0.3027 0.2081 0.4346 0.7674 1.4162 1.8693 2.5348
$8^{\circ}$ 0.6500 0.7995 0.9636 1.3000 1.5989 1.9272 0.1875 0.2725 0.3248 0.3000 0.5989 0.9272 1.6000 2.1979 2.8544

### v3

@@ Abstract
Note that the exponential functions in the numerator and denominator cancel each other out when calculating all considered ratios, while the scattering cross section itself rapidly decreases toward zero with increasing scattering angles at high energies, as previously reported in Refs. [14, 15]. Therefore, their plots remain fully visible across all scattering angles. However, our data table only lists the ratios at narrow angles since their values at larger angles are not particularly meaningful once all the scattering cross sections are extremely small. In general, the results in Table I indicate that the information obtained in each scattering process strongly depends on how electron polarization changes. Specifically, once electron polarization remains unchanged, the scattering cross section $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ is significantly larger than that obtained when the outgoing beam becomes unpolarized $\sigma_{\text{total}}^{\text{pol.}}(1,0)$, in particular at energies above 100 GeV and scattering angles $\theta=5^{\circ}-8^{\circ}$. Since the unpolarized term $\sigma_{\text{total}}^{0}$ in the total polarized cross section is identical in all cases, we can also make a direct comparison between $\sigma_{\text{total}}^{10}$ and $\sigma_{\text{total}}^{11}$. At energies below 10 GeV, the $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{11}$ ratios are very small, with the highest order of magnitude being only around $10^{-4}$ at $\theta=8^{\circ}$. However, at higher energy scales, these ratios increase remarkably, reaching up to several tens of percent. At the same energy and scattering angle, except for $\theta=0^{\circ}$, their magnitudes follow an ascending order of 7Li, 6Li, and 7Be nuclei. To further clarify the impact of electron polarization in each scattering process, let us now consider the ratios $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{0}$ and $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$. Similar to the $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{11}$ ratio, the contribution of $\sigma_{\text{total}}^{10}$ can be neglected at 10 GeV, since its highest order of magnitude relative to $\sigma_{\text{total}}^{0}$ at $\theta=8^{\circ}$ is only around $10^{-4}$. In contrast, $\sigma_{\text{total}}^{11}$ is nearly equal to $\sigma_{\text{total}}^{0}$ at energies not more than 10 GeV. At higher energy scales, the contribution of $\sigma_{\text{total}}^{10}$ becomes significant, whereas $\sigma_{\text{total}}^{11}$ even exceeds $\sigma_{\text{total}}^{0}$ by several times, as shown in Fig. 3 for 7Be. The unpolarized and polarized scattering cross sections in cm2/sr corresponding to several small scattering angles are also presented in Table II, considering the energy range that current electron accelerators can test. The results show that, at energies of several hundred GeV and above, the cross section approaches zero rapidly as the scattering angle increases. This implies that the electron could penetrate deeply inside the nucleus and nucleon, reaching the quark-size region and almost go straight after scattering, which is relevant to extremely small values of the cross section for nonzero scattering angles. The interaction between the electron and the nucleus is then more likely to be through smaller components, namely the quarks, rather than through its nucleons. In this sense, the scattering cross section ratios presented in Table I and Figs. 1-3 also provide important information that helps to clarify the role of electron polarization together with the weak interaction in the quark-size region, since the nucleon form factor in use decreases inversely with the square of the momentum transfer.
@@
Overall, all the ratios under consideration do not depend on the nuclear recoil factor and the center-of-mass motion. Furthermore, the calculations use the same harmonic oscillator parameter value for both polarized and unpolarized scattering, so that if we adjust the above parameters, the results are still the same. They show a little difference among selected nuclei in the scattering processes at energies below 10 GeV. At higher energies, their magnitudes increase in the order of 7Li, 6Li, and 7Be nuclei. Particularly, there are remarkable discrepancies in the contribution of electron polarization to the scattering cross section when comparing stable and unstable nuclei. The 7Li -7Be pair proves this remark, although both have the same nucleon numbers, in which 7Be is an unstable nucleus with a half-life of around 53.22 days. In contrast, for the 6Li -7Li stable pair, these discrepancies are much smaller, even though their nucleon numbers are slightly different. It can be explained that the weak interaction between electrons and nuclei strongly relates to nuclear stability. This property affects the correlation between the weak interaction and electron polarization, along with their contribution in each scattering process. Furthermore, although the $\sigma_{\text{total}}^{10}$ term can take negative values, its magnitude is significantly smaller than the $\sigma_{\text{total}}^{0}$ and $\sigma_{\text{total}}^{11}$ ones. Therefore, the total differential cross section also depends on the correlation between the electron spin orientations before and after scattering and remains positive in all cases. In particular, when the longitudinally polarized electrons move straight forward after scattering, the electron polarization-weak interaction correlation is not exhibited. This is evidenced by the ratios at $\theta=0^{\circ}$ in Table I. Their values remain nearly unchanged at all considered energies, except for $\sigma_{\text{total}}^{10}$, whose contribution can be omitted. So, besides nuclear spin orientation, it is evident that electron polarization and its transformation under the weak interaction have a significant impact on electron-nucleus scattering processes at high energies. The scattering cross section ratios in the cases $\sigma_{\text{total}}^{\text{pol.}}(-1,0)$ and $\sigma_{\text{total}}^{\text{pol.}}(-1,-1)$ can be analyzed similarly and we can extend to other light nuclei with the multipole form factors being determined by using the formulas in Ref. [15].
@@
In conclusion, we have performed a systematic study on the polarized electron scattering from unoriented light nuclei, utilizing the unified electroweak theory. Numerical calculations of the scattering cross sections in different electron polarization cases for three 6,7Li and 7Be light nuclei are obtained by directly calculating their multipole form factors. The results obtained are then compared and/or correlated with the corresponding cross sections for the unpolarized electron scattering case. It has been found that at the scattering angle of $\theta\simeq 0^{\circ}$, the longitudinal electron polarization does not explicitly correlate with the weak interaction for all energy scales. Such a correlation is strongly exhibited at the remaining scattering angles when the incoming electron energy passes 10 GeV. By comparing the results obtained for stable and unstable nuclei, we found that the nuclear stability may significantly influence the contribution of electron polarization to the scattering cross section. This research direction can be extended in many ways, specifically considering various state transitions within a single nucleus or different nuclei. Then, the asymmetry resulting solely from electron polarization and the weak interaction can be figured out. In addition, other interesting results are expected once the nuclear spin orientation and charge-changing weak current are included. Moreover, we could apply the multipole expansion for the scattering cross section within the unified electroweak theory to directly study the EMC effect [4, 25, 26, 27] in a target nucleus using both unpolarized and polarized electrons, e.g., through quasielastic electron scattering on 7Li and 7Be nuclei corresponding to a nuclear excitation from the ground state to the nearest state. More specifically, we shall calculate the cumulative cross section contributed by separate nucleons inside a certain nucleus and compare it with that made of a similar number of nucleons bound inside the same nucleus. This means that we can investigate the above effect on only a single nucleus instead of indirectly comparing its scattering cross section with that of a lighter nucleus, in which calculations are based on nucleons and nuclei instead of quarks. These extensive studies will provide further vital information, contributing to a thorough understanding of the nature of the EMC effect in inelastic electron-nucleus scattering.
@@
Table 1: Calculated cross section ratios for 6,7Li and 7Be nuclei obtained within the unified electroweak theory at different electron energies $\varepsilon$ = 10, 100, 500, and 1000 GeV and small scattering angles.
@@
10 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000
9.8236
$\times 10^{-10}$
4.6965
$\times 10^{-10}$
2.1329
$\times 10^{-10}$
9.8236
$\times 10^{-10}$
4.6965
$\times 10^{-10}$
2.1329
$\times 10^{-10}$
1.0000 1.0000 1.0000
$3^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000
2.1903
$\times 10^{-5}$
3.2235
$\times 10^{-5}$
4.8553
$\times 10^{-6}$
2.1903
$\times 10^{-5}$
3.2237
$\times 10^{-5}$
4.8554
$\times 10^{-6}$
1.0000 1.0001 1.0000
$5^{\circ}$ 0.5000 0.5000 0.5000 0.9999 1.0001 1.0000
-5.9973
$\times 10^{-5}$
8.9254
$\times 10^{-5}$
-3.4234
$\times 10^{-5}$
-5.9966
$\times 10^{-5}$
8.9270
$\times 10^{-5}$
-3.4231
$\times 10^{-5}$
0.9999 1.0002 0.9999
$8^{\circ}$ 0.4999 0.5001 0.4998 0.9998 1.0002 0.9996
-1.9758
$\times 10^{-4}$
2.2865
$\times 10^{-4}$
-3.5613
$\times 10^{-4}$
-1.9750
$\times 10^{-4}$
2.2876
$\times 10^{-4}$
-3.5588
$\times 10^{-4}$
0.9996 1.0005 0.9993
100 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000
9.8194
$\times 10^{-8}$
4.6965
$\times 10^{-8}$
2.1317
$\times 10^{-8}$
9.8194
$\times 10^{-8}$
4.6965
$\times 10^{-8}$
2.1317
$\times 10^{-8}$
1.0000 1.0000 1.0000
$3^{\circ}$ 0.5005 0.5016 0.5030 1.0010 1.0032 1.0059 0.0010 0.0032 0.0059 0.0010 0.0032 0.0059 1.0021 1.0064 1.0119
$5^{\circ}$ 0.5018 0.5044 0.5089 1.0036 1.0088 1.0179 0.0036 0.0087 0.0172 0.0036 0.0088 0.0179 1.0072 1.0177 1.0357
$8^{\circ}$ 0.5049 0.5112 0.5230 1.0097 1.0223 1.0461 0.0096 0.0214 0.0422 0.0097 0.0223 0.0461 1.0195 1.0447 1.0921
500 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000
2.4294
$\times 10^{-6}$
1.1741
$\times 10^{-6}$
5.2571
$\times 10^{-7}$
2.4294
$\times 10^{-6}$
1.1741
$\times 10^{-6}$
5.2571
$\times 10^{-7}$
1.0000 1.0000 1.0000
$3^{\circ}$ 0.5167 0.5371 0.5766 1.0334 1.0741 1.1533 0.0313 0.0645 0.1173 0.0334 0.0741 0.1533 1.0668 1.1482 1.3066
$5^{\circ}$ 0.5409 0.5894 0.6804 1.0818 1.1788 1.3609 0.0703 0.1317 0.2096 0.0818 0.1788 0.3609 1.1635 1.3577 1.7217
$8^{\circ}$ 0.5807 0.6722 0.8222 1.1614 1.3444 1.6444 0.1220 0.2039 0.2815 0.1614 0.3444 0.6444 1.3227 1.6889 2.2889
1000 $0^{\circ}$ 0.5000 0.5000 0.5000 1.0000 1.0000 1.0000
9.3906
$\times 10^{-6}$
4.6964
$\times 10^{-6}$
2.0144
$\times 10^{-6}$
9.3908
$\times 10^{-6}$
4.6965
$\times 10^{-6}$
2.0144
$\times 10^{-6}$
1.0000 1.0000 1.0000
$3^{\circ}$ 0.5545 0.6183 0.7335 1.1089 1.2365 1.4670 0.0894 0.1606 0.2415 0.1089 0.2365 0.4670 1.2179 1.4730 1.9340
$5^{\circ}$ 0.6037 0.7173 0.8839 1.2074 1.4346 1.7677 0.1466 0.2325 0.3028 0.2074 0.4346 0.7677 1.4148 1.8693 2.5355
$8^{\circ}$ 0.6498 0.7995 0.9635 1.2996 1.5989 1.9270 0.1873 0.2725 0.3248 0.2996 0.5989 0.9270 1.5992 2.1979 2.8541
Table 2: Numerically unpolarized and polarized electron elastic scattering cross sections (cm2/sr) for 6,7Li and 7Be nuclei obtained within the unified electroweak theory at different electron energies $\varepsilon$ = 10, 50, 100, and 120 GeV and small angles.
$\varepsilon$ $\theta$ 7Li 6Li 7Be
(GeV) $\sigma^{\text{unpol.}}_{\text{total}}$ $\sigma^{\text{pol.}}_{\text{total}}(1,0)$ $\sigma^{\text{pol.}}_{\text{total}}(1,1)$ $\sigma^{\text{unpol.}}_{\text{total}}$ $\sigma^{\text{pol.}}_{\text{total}}(1,0)$ $\sigma^{\text{pol.}}_{\text{total}}(1,1)$ $\sigma^{\text{unpol.}}_{\text{total}}$ $\sigma^{\text{pol.}}_{\text{total}}(1,0)$ $\sigma^{\text{pol.}}_{\text{total}}(1,1)$
10 $0^{\circ}$
5.1899
$\times 10^{-19}$
2.5949
$\times 10^{-19}$
5.1899
$\times 10^{-19}$
5.1898
$\times 10^{-19}$
2.5949
$\times 10^{-19}$
5.1898
$\times 10^{-19}$
2.0759
$\times 10^{-18}$
1.0379
$\times 10^{-18}$
2.0759
$\times 10^{-18}$
$3^{\circ}$
1.0666
$\times 10^{-30}$
0.5333
$\times 10^{-30}$
1.0666
$\times 10^{-30}$
6.3607
$\times 10^{-34}$
3.1803
$\times 10^{-34}$
6.3607
$\times 10^{-34}$
3.3791
$\times 10^{-30}$
1.6896
$\times 10^{-30}$
3.3791
$\times 10^{-30}$
$5^{\circ}$
9.7062
$\times 10^{-33}$
4.8531
$\times 10^{-33}$
9.7052
$\times 10^{-33}$
1.0157
$\times 10^{-33}$
0.5079
$\times 10^{-33}$
1.0158
$\times 10^{-33}$
1.9498
$\times 10^{-32}$
0.9749
$\times 10^{-32}$
1.9498
$\times 10^{-32}$
$8^{\circ}$
1.8223
$\times 10^{-37}$
0.9110
$\times 10^{-37}$
1.8219
$\times 10^{-37}$
1.7298
$\times 10^{-39}$
0.8651
$\times 10^{-39}$
1.7301
$\times 10^{-39}$
9.6850
$\times 10^{-38}$
4.8406
$\times 10^{-38}$
9.6811
$\times 10^{-38}$
50 $0^{\circ}$
2.0717
$\times 10^{-20}$
1.0358
$\times 10^{-20}$
2.0717
$\times 10^{-20}$
2.0713
$\times 10^{-20}$
1.0357
$\times 10^{-20}$
2.0713
$\times 10^{-20}$
8.2870
$\times 10^{-20}$
4.1435
$\times 10^{-20}$
8.2870
$\times 10^{-20}$
$2^{\circ}$
2.3128
$\times 10^{-40}$
1.1562
$\times 10^{-40}$
2.3123
$\times 10^{-40}$
1.1994
$\times 10^{-42}$
0.6005
$\times 10^{-42}$
1.2011
$\times 10^{-42}$
5.8293
$\times 10^{-41}$
2.9129
$\times 10^{-41}$
5.8252
$\times 10^{-41}$
$3^{\circ}$
9.0803
$\times 10^{-55}$
4.5401
$\times 10^{-55}$
9.0803
$\times 10^{-55}$
3.7204
$\times 10^{-58}$
1.8662
$\times 10^{-58}$
3.7323
$\times 10^{-58}$
1.4712
$\times 10^{-55}$
0.7357
$\times 10^{-55}$
1.4713
$\times 10^{-55}$
$5^{\circ}$
5.6798
$\times 10^{-100}$
2.8416
$\times 10^{-100}$
5.6832
$\times 10^{-100}$
1.0340
$\times 10^{-107}$
0.5215
$\times 10^{-107}$
1.0431
$\times 10^{-107}$
1.1469
$\times 10^{-100}$
0.5756
$\times 10^{-100}$
1.1513
$\times 10^{-100}$
100 $0^{\circ}$
5.1464
$\times 10^{-21}$
2.5732
$\times 10^{-21}$
5.1464
$\times 10^{-21}$
5.1422
$\times 10^{-21}$
2.5711
$\times 10^{-21}$
5.1422
$\times 10^{-21}$
2.0587
$\times 10^{-20}$
1.0294
$\times 10^{-20}$
2.0587
$\times 10^{-20}$
$1^{\circ}$
1.2618
$\times 10^{-39}$
0.6308
$\times 10^{-39}$
1.2615
$\times 10^{-39}$
6.6551
$\times 10^{-42}$
3.3276
$\times 10^{-42}$
6.6551
$\times 10^{-42}$
3.2308
$\times 10^{-40}$
1.6144
$\times 10^{-40}$
3.2285
$\times 10^{-40}$
$2^{\circ}$
2.7431
$\times 10^{-74}$
1.3718
$\times 10^{-74}$
2.7436
$\times 10^{-74}$
1.4419
$\times 10^{-79}$
0.7459
$\times 10^{-79}$
1.4918
$\times 10^{-79}$
5.0209
$\times 10^{-75}$
2.5150
$\times 10^{-75}$
5.0304
$\times 10^{-75}$
$3^{\circ}$
1.4273
$\times 10^{-130}$
0.7144
$\times 10^{-130}$
1.4287
$\times 10^{-130}$
2.0478
$\times 10^{-141}$
1.0999
$\times 10^{-141}$
2.1995
$\times 10^{-141}$
3.0663
$\times 10^{-131}$
1.5423
$\times 10^{-131}$
3.0844
$\times 10^{-131}$
120 $0^{\circ}$
3.5605
$\times 10^{-21}$
1.7803
$\times 10^{-21}$
3.5605
$\times 10^{-21}$
3.5564
$\times 10^{-21}$
1.7782
$\times 10^{-21}$
3.5564
$\times 10^{-21}$
1.4243
$\times 10^{-20}$
0.7121
$\times 10^{-20}$
1.4243
$\times 10^{-20}$
$1^{\circ}$
1.5070
$\times 10^{-44}$
0.7533
$\times 10^{-44}$
1.5067
$\times 10^{-44}$
3.7649
$\times 10^{-47}$
1.8825
$\times 10^{-47}$
3.7649
$\times 10^{-47}$
2.6684
$\times 10^{-45}$
1.3334
$\times 10^{-45}$
2.6665
$\times 10^{-45}$
$2^{\circ}$
3.9184
$\times 10^{-94}$
1.9604
$\times 10^{-94}$
3.9204
$\times 10^{-94}$
2.2680
$\times 10^{-101}$
1.2748
$\times 10^{-101}$
2.5497
$\times 10^{-101}$
7.7925
$\times 10^{-95}$
3.9095
$\times 10^{-95}$
7.8190
$\times 10^{-95}$
$3^{\circ}$
6.4150
$\times 10^{-175}$
3.2126
$\times 10^{-175}$
6.4259
$\times 10^{-175}$
3.2141
$\times 10^{-190}$
1.9873
$\times 10^{-190}$
3.9742
$\times 10^{-190}$
1.4418
$\times 10^{-175}$
0.7274
$\times 10^{-175}$
1.4548
$\times 10^{-175}$